# Axionlike particle-assisted supercooling chiral phase transition in QCD: Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenarios

### v2

@@ IV ALP assisted CW-type chiral phase transition in QCD
Another type of enhancement of the positive soft-scale breaking would be generated from the intrinsic $\mu_{5}$ corrections acting as the catalyzer of the chiral/axial breaking under a proper regularization scheme Farias:2016let . Including this assist, the size of $(y\,v_{\Phi})$ desired to realize the CW-type supercooling would be smaller. The detailed study to precisely achieve the ALP-assisted CW-type supercooling is noteworthy to be performed in another publication.
At sufficiently low temperatures much below the confinement scale of ${\cal O}(100\,{\rm MeV})$, including the expected nucleation at $T_{n}$, where the theory is already also in the chiral broken phase with $\langle{\cal S}\rangle\simeq f_{\pi}\simeq 92$ MeV (where the estimated value follows from input parameters in Eq.(15) with the estimated MF quark condensate $S\equiv\langle\bar{q}q\rangle\simeq(-330\,{\rm MeV})$). Then the $\mu_{5}$ terms will be gone because of decoupling of the QCD sphaleron, consequently, $(y\,v_{\Phi})$ also goes away. Still, however, the $\Phi$-ALP field gets negative-mass squared correction from the quark loop, enough to develop its vacuum expectation value in the confining and chiral broken phase. The thermodynamic potential in Eq.(19) indeed provides an effective potential for $\phi=(y\phi_{s})$ like
$\Omega(\phi=y\phi_{s})\Bigg|_{T\ll T_{c}}\sim-\frac{N_{c}N_{f}}{4\pi^{2}}\Lambda^{2}(f_{\pi}-\phi)^{2}+\frac{N_{f}N_{c}}{16\pi^{2}}(f_{\pi}-\phi)^{4}\ln\frac{\Lambda^{2}}{(f_{\pi}-\phi)^{2}}\,.$ (23)
@@
$(v_{\Phi})_{0}\approx-\left(\frac{y^{4}}{\lambda}\frac{N_{c}N_{f}}{2\pi^{2}}f_{\pi}\Lambda^{2}\right)^{1/3}\,$, (24)
for $(v_{\Phi})_{0}\ll f_{\pi}$. Here we have assumed $y^{4}\ll\lambda$, so that the $\lambda$-quartic term highly dominates over the loop-induced quartic term. This $(v_{\Phi})_{0}$ is generated essentially by the explicit breaking, via the $y$-Yukawa coupling term in Eq.(16) with $\langle\bar{q}q\rangle\neq 0$.
@@
$m_{\chi}\approx\frac{\sqrt{\lambda}}{y}|(v_{\Phi})_{0}|$ $\approx\left(\frac{\lambda^{1/2}}{y}\frac{N_{c}N_{f}}{2\pi^{2}}f_{\pi}\Lambda^{2}\right)^{1/3}$
$\simeq 214\,{\rm MeV}\times\left(\frac{\lambda^{1/2}}{y}\right)^{1/3}\,$, (26)
for $N_{c}=3$ and $N_{f}=2$. Here we have ignored the quark-loop induced wavefunction renormalization correction. The Nambu-Goldstone boson mode $a$ in $\Phi$ also gets the mass and the axionlike potential due to the generation of the explicit-breaking tadpole induced by $\langle\bar{q}q\rangle\neq 0$ in Eq.(16):
@@
$m_{a}^{2}=\frac{y\langle-\bar{q}q\rangle}{(-v_{\Phi})_{0}}$ $\simeq\frac{yf_{\pi}}{2(g_{s}+g_{d})(v_{\Phi})_{0}}$
$\simeq\left[124\,{\rm MeV}\times\left(\frac{\lambda}{y}\right)^{1/6}\right]^{2}\,$, (28)
where we have used the MFA $S\equiv\langle\bar{q}q\rangle$, which equals to ${\cal S}/(2(g_{s}+g_{d}))$ following from Eq.(20). Note that $m_{a}<m_{\chi}$ for any of $\lambda$ and $y$, as long as $y\ll 1$ and $\lambda={\cal O}(1)$.
@@
where $\Gamma_{f_{0}(500)/\eta^{\prime}}$ denotes the total widths of $f_{0}(500)$ and $\eta^{\prime}$, respectively. In Eq.(29) the denominator form comes from the $f_{0}(500)\eta^{\prime}$ propagator in the Breit-Wigner form, evaluated at the transfer momentum squared $q^{2}=m_{\chi/a}^{2}$, and the model-dependent factor $\sim y\Lambda^{2}$ denotes the mass mixing strength between the ALPs and those mesons. In the present evaluation, the latter part has been derived from the MFA-NJL based on the effective potential form in Eq.(19), with $\cal S$} and $v_{\Phi}$ replaced by $\eta^{\prime}$ and $(v_{\Phi})_{0}\sin a/(v_{\Phi})_{0}$ following from the polar decomposition of $\Phi$ in Eq.(25) applied to the $y$-Yukawa term in Eq.(16). The wavefunction renormalization of $f_{0}(500)/\eta^{\prime}$ has also been taken into account, which in the NJL-MFA yields $\sim N_{c}N_{f}/(16\pi^{2})\ln\Lambda^{2}/f_{\pi}^{2}$ for the inverse propagator in the chiral broken phase with $(v_{\Phi})_{0}\ll f_{\pi}$. Although the present NJL model only includes two flavors, the essential cutoff $\Lambda$ dependence on the mixing strength and wavefunction renormalization factor should be generic up to a factor of ${\cal O}(1)$, for the flavor singlet mesons, such as $\eta_{0}$ for the two-flavor case and $\eta^{\prime}_{0}$ for the three-flavor case. Therefore, the thus estimated net model-dependent factor $\sim y\Lambda^{2}$ in Eq.(29) covers the sufficient enough size in order of magnitude.
The $f_{0}(500)$ coupling to diphoton has not yet been clearly determined in experiments ParticleDataGroup:2024cfk . Moreover, the $\chi$ resonance will be contaminated with the $f_{0}(500)$ production, e.g., in the $\pi\pi$ scattering process. This channel will thus be challenging to use the probe of the present ALP-assisted model. Therefore, at this point, we disregard discussing the phenomenological constraint and consequence of $\chi$.
@@
Figure 3: The predicted curve of $g_{a\gamma\gamma}$ as a function of $m_{a}$ with $\lambda=1$ fixed. The plot range has taken into account the lower bound on $m_{a}$, set when $y=\lambda$ to ensure the perturbatively small enough quark loop correction to the ALP sector. The spike structure is generated at the resonance point when $m_{a}=m_{\eta^{\prime}}$, based on the resonant form factor structure of $g_{a\gamma\gamma}$ in Eq.(30). All the domains plotted in this figure survive over the current experimental constraint summarized in the webcite https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html.
@@
Solving Eq.(28) with respect to $y$; $y=(124\,{\rm MeV}/m_{a})^{6}\lambda$, and putting this $y$ into Eq.(30), we numerically write $g_{a\gamma\gamma}$ as a function of $m_{a}$. This function draws the theoretical prediction curve on the plane $(m_{a},g_{a\gamma\gamma})$ available at the AxionLimits (https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html). The predicted curve on $(m_{a},g_{a\gamma\gamma})$ is plotted in Fig. 3, where we have taken $\lambda=1$. We find that when $m_{a}\gtrsim 32\,{\rm MeV}$, where $g_{a\gamma\gamma}\lesssim 10^{-12}\,{\rm GeV}^{-1}$, the curve can be consistent with the current experiments, astrophysical, and cosmological observations. However, this lower bound on $m_{a}$ gives too large $y$: $y\sim\lambda$ in magnitude, therefore, the actual lower bound is set when $y=\lambda$, i.e., $m_{a}\simeq 124$ MeV. Figure 3 has already taken into account this limit. As a benchmark, we have $g_{a\gamma\gamma}\sim 10^{-17}$ for $m_{a}\sim 1$ GeV, which corresponds to the case with $y\sim 10^{-6}$ with $\lambda=1$ fixed.
This heavy ALP is unlikely to be probed or excluded even by the upcoming prospected experiments, such as the Electron Ion Collider experiment Balkin:2023gya . However, the indirect evidence could be tested via the GW and PBH productions sourced from the CW-type supercooling chiral phase transition, as has been suspected above. Detailed studies on this kind of cosmological observation feasibility are to be pursued in another publication.
@@ V Conclusion
In conclusion, we have clarified that the chiral phase transition in QCD-like theories can be CW type supercooling. This possibility paves a way to a new search window for new physics to be probed by GW and PBH productions. We have discussed this possibility by monitoring ordinary QCD setup in a view of a two-flavor NJL model in hot and/or dense medium, at the MFA level. It turned out that the identification essentially follows the scale violation classification: soft-scale breaking and quantum scale anomaly of the CW-type. The CW-type chiral-phase transition is possible to realize when ordinary QCD is coupled to or replaced by Beyond the Standard Model in a nontrivial way. A novel scenario includes prediction of an ALP with mass $\gtrsim$ sub GeV, which is likely to have a rich cosmology and phenomenological consequences at around the QCD scale, such as GW and PBH productions. More detailed study would be worth performing in the future.
The CW-type chiral phase transition is realized due to the dominance of the quantum scale anomaly of the logarithmic form at around the chiral criticality. This qualitative feature would still be operative even including gluonic or alternatively non-local (momentum dependent) interactions among quarks, because of persistence of the logarithmic renormalization group running for the potential (more precisely the induced quartic coupling of the CW scalar, which is identified as the bosonized quark bilinear in QCD-like theories). This characteristic logarithmic scale-anomaly is seen in any regularization scheme Kohyama:2015hix , unless dimensional reduction takes place as in the case with a strong magnetic field. Thus, our present proposal would be free from the intrinsic regularization dependence on the quantitative deformation of the chiral phase structure as has been addressed in the literature Kohyama:2015hix .

### v4

@@ IV.1 Cancellation of soft-scale breaking mass term in QCD plasma
Another type of enhancement of positive soft-scale breaking would be generated from intrinsic $\mu_{5}$ corrections acting as a catalyzer of the chiral/axial breaking under a proper regularization scheme Farias:2016let . Including this assist, the size of $(y\,v_{\Phi})$ in Eq.(32) would be smaller. The detailed study to more precisely achieve the ALP-assisted CW-type supercooling is noteworthy to be performed in another publication.
Figure 3: The plot of the thermal evolution of the effective potential $V_{\rm eff}$ in Eq.(27) with $M^{2}(\sigma)=0$ plus the Debye mass term correction in Eq.(33), normalized by the vacuum energy $V_{0}=|V_{\rm eff}(\langle\sigma_{0}\rangle)|$ measured at the true vacuum $\langle\sigma_{0}\rangle_{\rm true}\simeq 461$ MeV. The critical temperature of the first order phase transition is observed as $T_{c}\simeq 628$ MeV (corresponding to the moment denoted by red curve) The false vacuum is set away from the origin because of the $\mu_{5}$-induced tadpole term, which is at $\langle\sigma_{0}\rangle_{\rm false}\simeq 7.9$ MeV when $T=T_{c}$, as depicted in the zoomed-in figure on the bottom-left interior panel.
Simply adding to the effective potential in Eq.(27), with the $M^{2}=0$ condition in Eq.(31), the Debye mass terms arising from the quark (in Eq.(11)) and meson thermal fluctuations
$V_{T^{2}}=\left(\frac{N_{c}}{3}T^{2}+\frac{\lambda_{0}(\sigma_{0},0)}{4}T^{2}\right)\cdot\frac{\sigma_{0}^{2}}{2}\,$, (33)
we can roughly evaluate the thermal evolution of the total potential around the criticality. This rough prescription can work fine because it turns out that $T_{c}\gg\sigma_{0}$ around the true vacuum, so that higher orders in $(\sigma_{0}/T)$, including thermal cubic and CW-type log terms, can be subdominant. See Fig. 3, where the critical temperature is observed around $T\simeq 628$ MeV, at which temperature the true vacuum (at $\langle\sigma_{0}\rangle\simeq 461$ MeV) degenerates with the false vacuum (at $\langle\sigma_{0}\rangle\simeq 7.9$ MeV) created by the tadpole term $\propto\mu_{5}$. Note that the $\mu_{5}$-induced tadpole term is sizable, $C\simeq-(541\,{\rm MeV})^{3}$, which is much greater in magnitude than the current-quark mass-induced tadpole of ${\cal O}(m\cdot\Lambda^{2})={\cal O}(100\,{\rm MeV})$. As has been noted in the beginning of the present section (and also see footnote #5), all the $\mu_{5}$ term contributions will be dropped out when the QCD sphaleron gets confined and frozen out. In the present scenario, the confining phase is realized after passing through the supercooling, when $T=T_{n}\ll T_{c}\simeq 628$ MeV, which is expected to be at $T={\cal O}(100\,{\rm MeV})$ just like the ordinary QCD scenario.
The false vacuum still thermally shifts from $\langle\sigma_{0}\rangle_{\rm false}\simeq 7.9$ MeV at $T=T_{c}$ to a larger value as $T$ cools. At $T=T_{n}(\ll T_{c})\sim{\cal O}(100)$ MeV, the potential barrier is gone and the false vacuum is no longer an inflection point (i.e., gets unstable). At this moment, the $\sigma$ field starts to roll from the false vacuum down to the true vacuum at $\langle\sigma_{0}\rangle_{\rm true}\simeq 461$ MeV. This is a classical roll, not a quantum tunneling. The operated mechanism there is precisely the same as the dynamical realization of the CW-type small-field inflation Zhang:2023acu ; Liu:2024xrh . This classical roll includes several steps: a slow roll, which could cause a mini-inflation; a fast roll, which could be relevant to nonadiabatic particle and entropy productions; and an oscillation around the true vacuum. Thus, this potentially involves rich cosmology. GW and/or PHB production could be realized Zhang:2024vpp ; Zhang:2025kbu during the slow-roll (mini-inflation) and/or nonadiabatic fast roll (preheating, just like the QCD preheating Wang:2022ygk ; Wang:2024tda ), and/or oscillation (reheating). These intriguing issues deserve to be addressed in detail in another publication.
IV.2 ALP phenomenology in present-day-universe
Passing through the reheating process, the Universe cools down to sufficiently low temperatures much below the confinement scale of ${\cal O}(100\,{\rm MeV})$ (or the expected nucleation at $T_{n}$). At this point, the QCD sphaleron has already been decoupled because of the Boltzmann suppression for the sphaleron rate, so have been all the $\mu_{5}$ contributions in the effective potential, so that the large negative mass squared, i.e., the conventional dynamical symmetry breaking has been recovered. The meson loop corrections are then subleading – since even the pions act as heavy enough due to the negative mass squared – as also noted in the previous subsection, so the theory is settled back in the chiral broken phase simply following the original NJL-MF dynamics with the vacuum $\langle{\cal S}\rangle\simeq 394$ MeV. This is the vacuum characterizing the hadron phase in the present-day Universe. However, it turns out that the present-day Universe includes not only the ordinary QCD hadrons, but also light ALPs.
Although the $\mu_{5}$ terms are gone, the $\Phi$-ALP field still gets a mass squared correction from the quark loop, enough to develop its vacuum expectation value in the confining and chiral broken phase. Noting that the $\Phi$-ALP field does not couple to the meson sector up to the one-loop level, we focus only on the quarkonic sector to evaluate the development of the $\phi$ potential at this point. That is the thermodynamic potential form in Eq.(19) with $\langle{\cal S}\rangle$ being replaced by $\langle{\cal S}\rangle\simeq 394$ MeV, which indeed provides an effective potential for $\phi=(y\phi_{s})$ as follows:
$\Omega(\phi=y\phi_{s})\Bigg|_{T\ll T_{c}}\sim-\frac{N_{c}N_{f}}{4\pi^{2}}\Lambda^{2}(\langle{\cal S}\rangle-\phi)^{2}+\frac{N_{f}N_{c}}{16\pi^{2}}(\langle{\cal S}\rangle-\phi)^{4}\ln\frac{\Lambda^{2}}{(\langle{\cal S}\rangle-\phi)^{2}}\,.$ (34)
@@ IV.2 ALP phenomenology in present-day-universe
$(v_{\Phi})_{0}\approx-\left(\frac{y^{4}}{\lambda}\frac{N_{c}N_{f}}{2\pi^{2}}\langle{\cal S}\rangle\Lambda^{2}\right)^{1/3}\,$, (35)
for $(v_{\Phi})_{0}\ll\langle{\cal S}\rangle$. Here we have assumed $y^{4}\ll\lambda$, so that the $\lambda$-quartic term highly dominates over the loop-induced quartic term. This $(v_{\Phi})_{0}$ is generated essentially by the explicit breaking, via the $y$-Yukawa coupling term in Eq.(16) with $\langle\bar{q}q\rangle\neq 0$.
@@
$m_{\chi}\approx\frac{\sqrt{\lambda}}{y}|(v_{\Phi})_{0}|$ $\approx\left(\frac{\lambda^{1/2}}{y}\frac{N_{c}N_{f}}{2\pi^{2}}\langle{\cal S}\rangle\Lambda^{2}\right)^{1/3}$
$\simeq 346\,{\rm MeV}\times\left(\frac{\lambda^{1/2}}{y}\right)^{1/3}\simeq 318\,{\rm MeV}\times\left(\frac{\mu_{5}}{100\,{\rm MeV}}\right)^{1/3}\,$, (37)
for $N_{c}=3$ and $N_{f}=2$. Here we have ignored the quark-loop induced-wavefunction renormalization correction for $\Phi$, and in the last equality, we have used the parameter condition in Eq.(31). The Nambu-Goldstone boson mode $a$ in $\Phi$ also gets the mass and axionlike potential due to the generation of the explicit-breaking tadpole induced by $\langle\bar{q}q\rangle\neq 0$ in Eq.(16):
@@
$m_{a}^{2}$ $=\frac{y\langle-\bar{q}q\rangle}{(-v_{\Phi})_{0}}$
$\simeq\left[196\,{\rm MeV}\times\left(\frac{\lambda}{y}\right)^{1/6}\right]^{2}\simeq\left[26\,{\rm MeV}\times\left(\frac{y}{10^{-5}}\right)^{1/6}\cdot\left(\frac{\mu_{5}}{100\,{\rm MeV}}\right)^{1/3}\right]^{2}\,.$ (39)
Note that $m_{a}<m_{\chi}$ for any of $\lambda$ and $y$, as long as $y\ll 1$.
@@
where $\Gamma_{f_{0}(500)/\eta^{\prime}}$ denotes the total widths of $f_{0}(500)$ and $\eta^{\prime}$, respectively. In Eq.(40) the denominator form comes from the $f_{0}(500)\eta^{\prime}$ propagator in the Breit-Wigner form, evaluated at the transfer momentum squared $q^{2}=m_{\chi/a}^{2}$, and the model-dependent factor $\sim y\Lambda^{2}$ denotes the mass mixing strength between the ALPs and those mesons. In the present evaluation, the latter part has been derived from the NJL-MFA based on the effective potential form in Eq.(19), with $\cal S$} and $v_{\Phi}$ replaced by $\eta^{\prime}$ and $(v_{\Phi})_{0}\sin a/(v_{\Phi})_{0}$ following the polar decomposition of $\Phi$ in Eq.(36) applied to the $y$-Yukawa term in Eq.(16). The wavefunction renormalization of $f_{0}(500)/\eta^{\prime}$ has also been taken into account, which in the NJL-MFA yields $\sim\frac{N_{c}}{8\pi^{2}}\ln\frac{\Lambda^{2}}{\langle{\cal S}\rangle^{2}}$ as in Eq.(24) for the inverse propagator in the chiral broken phase with $(v_{\Phi})_{0}\ll\langle{\cal S}\rangle$. The thus estimated net model-dependent factor $\sim y\Lambda^{2}$ in Eq.(40) covers the sufficient size in order of magnitude.
The $f_{0}(500)$ coupling to diphoton has not yet been clearly determined in experiments ParticleDataGroup:2024cfk . Moreover, the $\chi$ resonance will be contaminated with the $f_{0}(500)$ production, e.g., in the $\pi\pi$ scattering process. This production channel will thus be challenging for the probe of the present ALP-assisted model. Therefore, at this point, we disregard discussing the phenomenological constraint and consequence of $\chi$.
@@
Figure 4: The predicted curve of $g_{a\gamma\gamma}$ as a function of $m_{a}$ with the soft-scale breaking-cancellation condition in Eq.(31) taken into account. The plot range ensures the perturbatively small enough quark loop correction to the ALP sector (i.e., $y^{4}\ll\lambda$, or equivalently $m_{a}\ll 169$ MeV). All the domains plotted not overlapping with the constraint regimes survive over the current experimental constraint summarized in the website https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html. The case with $m_{a}<1$ MeV has also been ruled out by the bound from the “Cosmic Background”.
@@
Solving Eq.(39) with respect to $y$; $y\simeq 10^{-5}\times(m_{a}/26\,{\rm MeV})^{6}$ for $\mu_{5}=100$ MeV, and putting this $y$ into Eq.(41), we numerically write $g_{a\gamma\gamma}$ as a function of $m_{a}$. This function draws the theoretical prediction curve on the plane $(m_{a},g_{a\gamma\gamma})$ available at the AxionLimits (https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html). The predicted curve on $(m_{a},g_{a\gamma\gamma})$ is plotted in Fig. 4, with the soft-scale breaking-cancellation condition in Eq.(31) and the perturbativity $y^{4}\lambda$ taken into account (i.e., $y^{4}\ll\lambda$, or equivalently $m_{a}\ll 169$ MeV). The $\mu_{5}$ sensitivity on $g_{a\gamma\gamma}$ is only quadratic: $g_{\rm a\gamma\gamma}\propto\mu^{2}_{5}$, so the predictive curve in Fig. 4 does not substantially change as long as $\mu_{5}$ is around the typical QCD scale, in a range of 100 - 200 MeV.
The $a$-ALP cannot be heavier than $\sim 169$ MeV so as to keep the perturbatively small enough quark loop correction to the ALP sector (i.e., $y^{4}\ll\lambda$). Furthermore, the current bound severely constrains $m_{a}$ to be $\sim 4$ MeV - 8 MeV or $\sim 1$ - 2.5 MeV, which has been reflected in the figure. The case with $m_{a}<1$ MeV has also been ruled out by the bound from the “Cosmic Background”. As a benchmark, we have $g_{a\gamma\gamma}\sim 10^{-12}\,{\rm GeV}^{-1}$ for $m_{a}\sim 5$ MeV, which corresponds to the case with $y\sim 5\times 10^{-10}$ with $\lambda\sim 10^{-19}$.
This heavy ALP is unlikely to be probed or excluded by the upcoming prospect experiments, such as the Electron Ion Collider experiment Balkin:2023gya or the Search for Hidden Particles (SHiP, at CERN) experiment and the Beam Dump eXperiment (BDX, at JLab) Patrone:2025fwk . However, indirect evidence could be tested via the GW and PBH productions sourced from the CW-type supercooling chiral phase transition, as has been suspected in the previous section.
@@ V Conclusion
In conclusion, we have clarified that the chiral phase transition in QCD-like theories can be CW type supercooling. This possibility paves the way to a new search window for new physics to be probed by GW and PBH productions. We have discussed this possibility by monitoring ordinary QCD setup in a view of a two-flavor NJL model in a hot and/or dense medium, at the MFA level. It turned out that the identification essentially follows the scale violation classification: soft-scale breaking and quantum scale anomaly of the CW-type. The CW-type chiral-phase transition is possible to realize when ordinary QCD is coupled to or replaced by Beyond the Standard Model in a nontrivial way. This is what is called the ALP-assisted CW-type supercooled chiral phase transition. This new scenario predicts rich cosmological and phenomenological consequences around the QCD scale: a post supercooled CW-type small-field (mini-) inflation; preheating; reheating along with GW and/or PBH production.
The proposed scenario predicts a heavy ALP with mass $\sim$ 5 MeV in the present-day Universe. Since the couplings $y$ and $\lambda$ are tiny, this ALP could also resolve the strong CP problem in a way similar to the conventional axion solution, consistently with the current bound from the neutron electric dipole moment. The natural inflation of this ALP can also be realized in the earlier epoch of the thermal history. The cosmological evolution from the inflationary to QCD epochs would also be worth pursuing.
@@
In closing, the CW-type chiral phase transition is realized because of the dominance of the quantum scale anomaly of the logarithmic form around the chiral criticality. This qualitative feature would still be operative even including gluonic or alternatively non-local (momentum dependent) interactions among quarks, because of persistence of the logarithmic renormalization group running for the potential (more precisely the induced quartic coupling of the CW scalar, which is identified as the bosonized quark bilinear in QCD-like theories). This characteristic logarithmic scale-anomaly is seen in any regularization scheme Kohyama:2015hix , unless dimensional reduction takes place as in the case with a strong magnetic field. Thus, our present proposal would be free from the intrinsic regularization dependence on the quantitative deformation of the chiral phase structure as has been addressed in the literature Kohyama:2015hix .